# Perpendicular ion heating in turbulence and reconnection: magnetic moment breaking by coherent fluctuations

### v1

@@ 1 Introduction
What is the source of free energy for the observed heating? One successful model is heating from the Alfvénic plasma turbulence ubiquitous in the solar wind (Belcher & Davis, 1971; Chen, 2016; Chen et al., 2020) and corona (De Pontieu et al., 2007): the fluctuation amplitudes are consistent with the observed plasma heating (Chandran & Hollweg, 2009; Cranmer et al., 2009), suggesting that the solar wind is accelerated and locally heated by the dissipation of these turbulent fluctuations. It is worth noting that, because the turbulence has only a very small compressive component (Klein et al., 2012), the ion heating within the gyrokinetic approximation would be much small to explain the observations (Schekochihin et al., 2009, 2019; Kawazura et al., 2020).
@@
Several theoretical models have been proposed to explain ion heating in turbulence. First, cyclotron resonant heating (Hollweg & Isenberg, 2002; Chandran et al., 2010; Isenberg & Vasquez, 2011, 2019; Bowen et al., 2024) occurs when, in the frame moving with an ion’s parallel velocity $v_{\parallel}$, a wave’s frequency $\omega-k_{\parallel}v_{\parallel}$ matches the gyrofrequency $\Omega_{i}$: hence "resonant", $\omega-k_{\parallel}v_{\parallel}-n\Omega_{i}=0$. This is often discussed in the framework of quasilinear theory (Kennel & Engelmann, 1966; Stix, 1992), where the resulting diffusion of energy in phase space is derived assuming a spectrum of infinite plane waves and considering only the resonant response.
@@
Another important model, closer in approach to that of the present paper, is stochastic heating (McChesney et al., 1987; Chandran et al., 2010), in which ions random-walk in energy due to uncorrelated kicks from ion-scale fluctuations. In the Chandran et al. (2010) model, this results in a heating rate $Q_{\perp}\sim\delta u_{\rho}^{3}/\rho_{\rm th}\exp(-c_{2}v_{\rm th}/\delta u_{\rho})$, where $\delta u_{\rho}$ is the amplitude of $E\times B$ velocity fluctuations at the gyroscale $\rho_{\rm th}$, and the exponential suppression factor was added empirically to account for the near-conservation of the magnetic moment at low frequencies and small amplitudes. One advantage of stochastic heating as opposed to cyclotron resonant heating is that it does not require an exact resonance, nor does it assume that the fluctuations resemble infinite plane waves. This allows one to easily incorporate the observed intermittent probability distribution of fluctuation amplitudes (Chandran et al., 2015; Mallet & Schekochihin, 2017) at the gyroscale, which can dramatically increase the predicted heating rate (Mallet et al., 2019; Cerri et al., 2021).
@@
Observations (Chen et al., 2011) show that the solar wind turbulence is highly anisotropic: fluctuations have very different characteristic lengthscales parallel ($l_{\parallel}$) and perpendicular ($\lambda$) to the background magnetic field, $l_{\parallel}\gg l_{\perp}$. Modern turbulence theories (Goldreich & Sridhar, 1995; Boldyrev, 2006) explain this in terms of a critical balance between characteristic timescales associated with linear propagation ($\tau_{lin}\sim l_{\parallel}/v_{\mathrm{A}}$, with $v_{\mathrm{A}}=B_{0}/\sqrt{4\pi n_{p}m_{p}}$ the Alfvén velocity based on the mean magnetic field $B_{0}$) and nonlinear interactions ($\tau_{nl}\sim\lambda/\delta u_{\lambda}$): the cascade time $\tau\sim\tau_{lin}\sim\tau_{nl}$, whence $l_{\parallel}/\lambda\sim v_{\mathrm{A}}/\delta u_{\lambda}\gg 1$. Since the fluctuating field amplitude $\delta u_{\lambda}$ at scale $\lambda$ is an increasing function of $\lambda$, at progressively smaller scales, the anisotropy $l_{\parallel}/\lambda$ increases. This means that the fluctuations remain relatively low-frequency, with $\omega\sim 1/\tau\ll\Omega_{i}$, where $\Omega_{i}=Z_{i}eB/m_{i}c$ is the ion gyrofrequency. This poses a challenge: the magnetic moment $\mu=m_{i}v_{\perp}^{2}/B$ is conserved to all orders in $\eta\sim\omega/\Omega_{i}\ll 1$ (Kruskal, 1962), and so the usual perturbation theory would suggest that perpendicular ion heating should be irrelevant for such anisotropic, small-amplitude turbulence, in contrast to the observations. It is worth noting that "to all orders" is not the same as "exactly": as an example (that will be important in this paper), $\exp(-1/\eta)\neq 0$, but is "zero to all orders" if $\eta\ll 1$, since all derivatives vanish as $\eta\to 0$.
@@
Besides turbulence, magnetic reconnection has been proposed as a mechanism for coronal heating (Klimchuk, 2015) and also as a heating mechanism within the turbulence itself (Shay et al., 2018). In fact, turbulent heating and reconnection heating may not be as distinct as traditionally thought. The turbulent cascade naturally leads to the formation of extended current sheets (Boldyrev, 2006; Chandran et al., 2015; Mallet & Schekochihin, 2017), which reconnect once their width becomes sufficiently small (Mallet et al., 2017, 2017; Boldyrev & Loureiro, 2017; Loureiro & Boldyrev, 2017a, b; Vech et al., 2018; Comisso et al., 2017; Cerri & Califano, 2017; Franci et al., 2017; Dong et al., 2022). Similarly, approaching the problem from the "reconnection end", extended reconnecting current sheets are often violently unstable, leading naturally to strong turbulence (Loureiro et al., 2007; Bhattacharjee et al., 2009; Huang & Bhattacharjee, 2016). Seeking to explain preferential heavy ion heating in the corona and solar flares, Drake et al. (2009a) developed a theory of perpendicular ion heating in reconnection exhausts, supported by numerical simulations. In Drake et al. (2009b), they showed that in guide-field reconnection, strong ion heating only occurs if the characteristic timescale to transit the exhaust is shorter than the ion’s gyroperiod. This behaviour has similarities to the stochastic heating in turbulence.
@@
All three of cyclotron-resonant, stochastic, and reconnection perpendicular ion heating share a common feature: they require the conservation of the magnetic moment to be broken. In the case of stochastic and reconnection heating, this leads to a "threshold" which must be satisfied for strong ion heating to be possible. Likewise, in cyclotron resonant heating, the resonance condition must be satisfied (we will argue that this "sharper" behaviour is a consequence of the plane-wave assumption). Johnston et al. (2025) noticed the similarities between cyclotron-resonant and stochastic heating, and found that the heating in their test-particle simulations was well described by a single exponential suppression factor modelling both cyclotron-resonant heating in imbalanced turbulence and stochastic heating in balanced turbulence.
@@
In this paper, we develop a new framework that describes perpendicular ion heating. We analytically study the response of an ion to a localized, coherent fluctuation in the electromagnetic fields, with the fluctuating electric $\delta\boldsymbol{E}$ and magnetic $\delta\boldsymbol{B}$ fields tending to zero at $t=\pm\infty$, an approach to our knowledge first taken in Krall & Rosenbluth (1964) and for general adiabatic invariants in Landau & Lifshitz (1976). Quite generically, we find that the perpendicular ion kinetic energy $m_{i}v_{\perp}^{2}/2$ changes by an amount of order
@@ 3 Solution for $\varepsilon\ll\eta_{K}\sim 1$
and proceed with our calculation. [^1] As mentioned above, in this section we formally require that the fields do not vary too slowly at small scales: $\varepsilon\ll\eta_{K}\sim 1$ for all $K\gtrsim 1$: this does not preclude $\eta_{K}\ll 1$, so long as $\eta_{K}\gg\varepsilon$. We will point out clearly where the calculation breaks down for the case of $\eta_{K}\sim\varepsilon$: and in Appendix C, we will show that this can be ”fixed”, with no change to our final results. At zeroth order in $\varepsilon$, we just have the gyration of the particle about the background field,
@@
$\dot{Y}_{1}=\int_{-\infty}^{T}\cos(T-T^{\prime})g(\sin T^{\prime},0,T^{\prime})\mathrm{d}\tau^{\prime}.$ (18)
@@
$\dot{Y}_{1}$ $=\frac{1}{2\pi}\cos T\int_{-\infty}^{T}\int_{-\infty}^{\infty}g(K,0,\eta_{K}T^{\prime})\sum_{n=-\infty}^{\infty}\frac{nJ_{n}(K)}{K}e^{inT^{\prime}}\mathrm{d}K\mathrm{d}T^{\prime}$
$\quad\,+\frac{1}{2\pi}\sin T\int_{-\infty}^{T}\int_{-\infty}^{\infty}g(K,0,\eta_{K}T^{\prime})\sum_{n=-\infty}^{\infty}\frac{J_{n-1}(K)-J_{n+1}(K)}{2i}e^{inT^{\prime}}\mathrm{d}K\mathrm{d}T^{\prime}.$ (22)
@@
$\dot{X}_{1}$ $=\frac{1}{2\pi}\sin T\int_{-\infty}^{T}\int_{-\infty}^{\infty}g(K,0,\eta_{K}T^{\prime})\sum_{n=-\infty}^{\infty}\frac{nJ_{n}(K)}{K}e^{inT^{\prime}}\mathrm{d}K\mathrm{d}T^{\prime}$
$\quad\,-\frac{1}{2\pi}\cos T\int_{-\infty}^{T}\int_{-\infty}^{\infty}g(K,0,\eta_{K}T^{\prime})\sum_{n=-\infty}^{\infty}\frac{J_{n-1}(K)-J_{n+1}(K)}{2i}e^{inT^{\prime}}\mathrm{d}K\mathrm{d}T^{\prime}.$ (23)
@@ 3.1 Change in perpendicular energy
$\frac{\langle v_{\perp}^{2}\rangle}{v_{\perp 0}^{2}}=1+2\varepsilon\left(\dot{X}_{0}\dot{X}_{1}+\dot{Y}_{0}\dot{Y_{1}}\right)_{T\to\infty}+O(\varepsilon^{2}).$ (25)
@@
$\Delta\sim 2i\int_{-\infty}^{\infty}\sum_{n=-\infty}^{\infty}\textrm{sgn}(n)\,\textrm{Res}[A(n,K,s),s_{*}]\exp\left(-\frac{|n\mathrm{Im}\left\{s_{*}\right\}|}{\eta(K)}\right)\mathrm{d}K.$ (28)
@@ 3.3 Scattering contours
Let us for the moment assume that the electromagnetic fluctuations are from a propagating wave or superposition of waves, with a parallel phase velocity $v_{\mathrm{ph}}(K)$, so that $\tilde{b}=\tilde{b}(K,Z-(v_{\mathrm{ph}}(K)/v_{\perp 0})T)$, and $\tilde{g}=\tilde{g}(K,Z-(v_{\mathrm{ph}}/v_{\perp 0})T)$. The results derived in the previous sections do not require this, but it will allow us to make contact with the usual quasilinear theory of cyclotron heating. It may also be directly applicable to the turbulence in the solar wind, which can be highly imbalanced: dominated by outward-going Alfvén waves, for which $v_{\mathrm{ph}}=v_{\mathrm{A}}$. More generally, this situation could potentially also apply to nonlinear solitary waves which have a single effective $v_{\mathrm{ph}}$ due to the nonlinearity balancing dispersion (Kakutani & Ono, 1969; Kawahara, 1969; Hasegawa & Mima, 1976; Mjølhus & Wyller, 1986; Mallet, 2023). We note that we are ignoring the possibility of purely waves with phase velocity only in the perpendicular direction, and also electrostatic waves (which to have parallel phase velocities must also have $E_{z}\neq 0$ so that Faraday’s law is satified).
@@ 4 Low-$\beta$ Alfvénic turbulence
In the Alfvénic turbulence present in the solar wind and corona, both $\varepsilon$ and $\eta$ depend on $k_{\perp}$. We will assume a relatively low $\beta$, so that the kinetic reduced electron heating model (KREHM) equations (Zocco & Schekochihin, 2011) may be used; we also assume that while $k_{\perp}\rho_{p}\sim k_{\perp}\rho_{s}\sim 1$, $k_{\perp}d_{e}\ll 1$, so the electrons are isothermal; where the thermal proton gyroradius $\rho_{p}=v_{th\rm{p}}/\Omega_{p}$ (different from $\rho$!), the ion sound radius is $\rho_{s}=\sqrt{ZT_{e}/m_{i}}/\Omega_{p}$, with $\Omega_{p}=eB/m_{p}c$ the proton gyrofrequency, and $d_{e}=c/\omega_{pe}$ is the electron inertial length. We want to express our results in terms of what is experimentally observable in the solar wind; namely, the $\delta B_{\perp}$ fluctuation amplitude as a function of the perpendicular wavenumber $k_{\perp}$; we will write this in velocity units as $\delta b_{k}=\delta B_{k}/\sqrt{4\pi n_{0i}m_{i}}$. For now, we will also neglect intermittency in the turbulent fluctuation amplitude, supposing that we may characterise the amplitude at each scale by a single value $\delta b_{k}$. (The effects of intermittency will be examined in Sec. 4.4.) Moreover, we assume the critical balance, such that $\omega\sim k_{\perp}\delta u_{e}$, where the effective electron bulk flow velocity is (Zocco & Schekochihin, 2011)
@@
This is true not only for linear Alfvén waves, but statistically even in strongly nonlinear kinetic-Alfvén turbulence (Grošelj et al., 2018): in a similar sense to the fact that $\delta u\sim\delta b$ in strong MHD turbulence (Maron & Goldreich, 2001). Since $\Gamma_{0}\approx 1-k_{\perp}^{2}\rho_{p}^{2}/2$ for $k_{\perp}\rho_{p}\ll 1$, but $\Gamma_{0}\approx 0$ for $k_{\perp}\rho_{p}\gg 1$, $\alpha_{k}\to 1$ as $k_{\perp}\rho_{p}\ll 1$ (Alfvén waves) and $\alpha_{k}\to k_{\perp}\rho_{\rm th}$ for $k_{\perp}\rho_{p}\gg 1$ (kinetic Alfvén waves). Thus, we have
@@
where we have neglected $k_{\parallel}v_{\parallel 0}\ll\omega\sim k_{\parallel}v_{\mathrm{A}}$ since $\beta_{i}$ is small, and as promised added an undetermined constant $c_{2}$ accounting for (several) prefactors of order unity we have neglected. This could in principle be corrected for the slowing down of the turbulent cascade due to dynamic alignment (Boldyrev, 2006; Chandran et al., 2015; Mallet & Schekochihin, 2017) and/or imbalance (Schekochihin, 2022; Chandran et al., 2025). A more important limitation is that the KREHM equations assume that $\omega/\Omega_{p}\ll 1$. We will rather flagrantly ignore this restriction in the following analysis: while it could be corrected for by using a more accurate dispersion relation, we believe it does not impact our results in a significant way. The parameter $\varepsilon_{k_{\perp}}=Ec/Bv_{\perp 0}$ controlling the amplitude of the electric fields is similarly found: using $E\sim k_{\perp}\phi$ and using Eqs. (49) and (51),
@@
For the moment assuming that the ion component of the plasma is mainly protons, $v_{\rm th}=v_{th\rm p}$, $\rho_{\rm th}\sim\rho_{p}$, and $\Omega_{i}=\Omega_{p}$, we have
@@
We will return to the subject of minor ions in Sec. 4.5. Taking $k\rho_{p}\sim 1$, we recover the expression given in Chandran et al. (2010). However, we can now assess the scale dependence of the heating directly. It is interesting to look at this in different limits. For $k_{\perp}\rho_{p}\ll 1$, we have $1-\Gamma_{0}\approx k_{\perp}^{2}\rho_{p}^{2}/2$, $\alpha_{k}\sim 1$, and $J_{1}(k_{\perp}\rho_{p})\sim k_{\perp}\rho_{p}/2$. Then,
@@
The only difference is in the exponential suppression factor; the speedup in the frequency for small-scale fluctuations means we get an extra factor of $k_{\perp}\rho_{p}$ there. Were $\delta b_{k}$ independent of scale, the heating rate becomes monotonically larger towards smaller scale, since the frequency increases.

### v2

@@ 1 Introduction
What is the source of free energy for the observed heating? One successful model is heating from the Alfvénic plasma turbulence ubiquitous in the solar wind (Belcher and Davis, 1971; Chen, 2016; Chen et al., 2020) and corona (De Pontieu et al., 2007): the fluctuation amplitudes are consistent with the observed plasma heating (Chandran and Hollweg, 2009; Cranmer et al., 2009), suggesting that the solar wind is accelerated and locally heated by the dissipation of these turbulent fluctuations. It is worth noting that, because the turbulence has only a very small compressive component (Klein et al., 2012), the ion heating within the gyrokinetic approximation would be much small to explain the observations (Schekochihin et al., 2009, 2019; Kawazura et al., 2020).
@@
Several theoretical models have been proposed to explain ion heating in turbulence. First, cyclotron resonant heating (Hollweg and Isenberg, 2002; Chandran et al., 2010; Isenberg and Vasquez, 2011, 2019; Bowen et al., 2024) occurs when, in the frame moving with an ion’s parallel velocity $v_{\parallel}$, a wave’s frequency $\omega-k_{\parallel}v_{\parallel}$ matches the gyrofrequency $\Omega_{i}$: hence "resonant", $\omega-k_{\parallel}v_{\parallel}-n\Omega_{i}=0$. This is often discussed in the framework of quasilinear theory (Kennel and Engelmann, 1966; Stix, 1992), where the resulting diffusion of energy in phase space is derived assuming a spectrum of infinite plane waves and considering only the resonant response.
@@
Another important model, closer in approach to that of the present paper, is stochastic heating (McChesney et al., 1987; Chandran et al., 2010), in which ions random-walk in energy due to uncorrelated kicks from ion-scale fluctuations. In the Chandran et al. (2010) model, this results in a heating rate $Q_{\perp}\sim\delta u_{\rho}^{3}/\rho_{\rm th}\exp(-c_{2}v_{\rm th}/\delta u_{\rho})$, where $\delta u_{\rho}$ is the amplitude of $E\times B$ velocity fluctuations at the gyroscale $\rho_{\rm th}$, and the exponential suppression factor was added empirically to account for the near-conservation of the magnetic moment at low frequencies and small amplitudes. One advantage of stochastic heating as opposed to cyclotron resonant heating is that it does not require an exact resonance, nor does it assume that the fluctuations resemble infinite plane waves. This allows one to easily incorporate the observed intermittent probability distribution of fluctuation amplitudes (Chandran et al., 2015; Mallet and Schekochihin, 2017) at the gyroscale, which can dramatically increase the predicted heating rate (Mallet et al., 2019; Cerri et al., 2021).
@@
Observations (Chen et al., 2011) show that the solar wind turbulence is highly anisotropic: fluctuations have very different characteristic lengthscales parallel ($l_{\parallel}$) and perpendicular ($\lambda$) to the background magnetic field, $l_{\parallel}\gg l_{\perp}$. Modern turbulence theories (Goldreich and Sridhar, 1995; Boldyrev, 2006) explain this in terms of a critical balance between characteristic timescales associated with linear propagation ($\tau_{lin}\sim l_{\parallel}/v_{\mathrm{A}}$, with $v_{\mathrm{A}}=B_{0}/\sqrt{4\pi n_{p}m_{p}}$ the Alfvén velocity based on the mean magnetic field $B_{0}$) and nonlinear interactions ($\tau_{nl}\sim\lambda/\delta u_{\lambda}$): the cascade time $\tau\sim\tau_{lin}\sim\tau_{nl}$, whence $l_{\parallel}/\lambda\sim v_{\mathrm{A}}/\delta u_{\lambda}\gg 1$. Since the fluctuating field amplitude $\delta u_{\lambda}$ at scale $\lambda$ is an increasing function of $\lambda$, at progressively smaller scales, the anisotropy $l_{\parallel}/\lambda$ increases. This means that the fluctuations remain relatively low-frequency, with $\omega\sim 1/\tau\ll\Omega_{i}$, where $\Omega_{i}=Z_{i}eB/m_{i}c$ is the ion gyrofrequency. This poses a challenge: the magnetic moment $\mu=m_{i}v_{\perp}^{2}/B$ is conserved to all orders in $\eta\sim\omega/\Omega_{i}\ll 1$ (Kruskal, 1962), and so the usual perturbation theory would suggest that perpendicular ion heating should be irrelevant for such anisotropic, small-amplitude turbulence, in contrast to the observations. It is worth noting that "to all orders" is not the same as "exactly": as an example (that will be important in this paper), $\exp(-1/\eta)\neq 0$, but is "zero to all orders" if $\eta\ll 1$, since all derivatives vanish as $\eta\to 0$.
@@
Besides turbulence, magnetic reconnection has been proposed as a mechanism for coronal heating (Klimchuk, 2015) and also as a heating mechanism within the turbulence itself (Shay et al., 2018). In fact, turbulent heating and reconnection heating may not be as distinct as traditionally thought. The turbulent cascade naturally leads to the formation of extended current sheets (Boldyrev, 2006; Chandran et al., 2015; Mallet and Schekochihin, 2017), which reconnect once their width becomes sufficiently small (Mallet et al., 2017a, b; Boldyrev and Loureiro, 2017; Loureiro and Boldyrev, 2017b, a; Vech et al., 2018; Comisso et al., 2017; Cerri and Califano, 2017; Franci et al., 2017; Dong et al., 2022). Similarly, approaching the problem from the "reconnection end", extended reconnecting current sheets are often violently unstable, leading naturally to strong turbulence (Loureiro et al., 2007; Bhattacharjee et al., 2009; Huang and Bhattacharjee, 2016). Seeking to explain preferential heavy ion heating in the corona and solar flares, Drake et al. (2009b) developed a theory of perpendicular ion heating in reconnection exhausts, supported by numerical simulations. In Drake et al. (2009a), they showed that in guide-field reconnection, strong ion heating only occurs if the characteristic timescale to transit the exhaust is shorter than the ion’s gyroperiod. This behaviour has similarities to the stochastic heating in turbulence.
@@
All three of cyclotron-resonant, stochastic, and reconnection perpendicular ion heating share a common feature: they require the conservation of the magnetic moment to be broken. In the case of stochastic and reconnection heating, this leads to a "threshold" which must be satisfied for strong ion heating to be possible. Likewise, in cyclotron resonant heating, the resonance condition must be satisfied (we will argue that this "sharper" behaviour is a consequence of the plane-wave assumption). Johnston et al. (2025) noticed the similarities between cyclotron-resonant and stochastic heating, and found that the heating in their test-particle simulations was well described by a single exponential suppression factor modelling both cyclotron-resonant heating in imbalanced turbulence and stochastic heating in balanced turbulence. In related work in a different physical setting, namely electron scattering in the radiation belts, a similarly close relationship between electron scattering by strongly nonlinear coherent structures (electron holes) and quasilinear theory was also previously discussed by Vasko et al. (2017) and Vasko et al. (2018).
@@
In this paper, we develop a new framework that describes perpendicular ion heating. We analytically study the response of an ion to a localized, coherent fluctuation in the electromagnetic fields, with the fluctuating electric $\delta\boldsymbol{E}$ and magnetic $\delta\boldsymbol{B}$ fields tending to zero at $t=\pm\infty$, an approach to our knowledge first taken in Krall and Rosenbluth (1964) for electric field fluctuations and for general adiabatic invariants in Landau and Lifshitz (1976). Quite generically, we find that the perpendicular ion kinetic energy $m_{i}v_{\perp}^{2}/2$ changes by an amount of order
@@ 3 Solution for $\varepsilon\ll\eta_{K}\sim 1$
and proceed with our calculation. [^1] As mentioned above, in this section we formally require that the fields do not vary too slowly at small scales: $\varepsilon\ll\eta_{K}\sim 1$ for all $K\gtrsim 1$: this does not preclude $\eta_{K}\ll 1$, so long as $\eta_{K}\gg\varepsilon$. We will point out clearly where the calculation breaks down for the case of $\eta_{K}\sim\varepsilon\ll 1$: and in Appendix C, we will show that this can be ”fixed”, with no change to our final results. Note also that our orderings $\varepsilon\ll\eta_{K}\sim 1$ and $K\sim 1$ are different from the usual guiding center regime, for which $\eta_{K}\ll 1$, $K\ll 1$, and $\epsilon\sim 1$. At zeroth order in $\varepsilon$, we just have the gyration of the particle about the background field,
@@
$\dot{Y}_{1}=\int_{-\infty}^{T}\cos(T-T^{\prime})g(\sin T^{\prime},0,T^{\prime})\mathrm{d}T^{\prime}.$ (18)
@@
$\dot{Y}_{1}$ $=\frac{1}{2\pi}\cos T\int_{-\infty}^{T}\int_{-\infty}^{\infty}\tilde{g}(K,0,\eta_{K}T^{\prime})\sum_{n=-\infty}^{\infty}\frac{nJ_{n}(K)}{K}e^{inT^{\prime}}\mathrm{d}K\mathrm{d}T^{\prime}$
$\quad\,+\frac{1}{2\pi}\sin T\int_{-\infty}^{T}\int_{-\infty}^{\infty}\tilde{g}(K,0,\eta_{K}T^{\prime})\sum_{n=-\infty}^{\infty}\frac{J_{n-1}(K)-J_{n+1}(K)}{2i}e^{inT^{\prime}}\mathrm{d}K\mathrm{d}T^{\prime}.$ (22)
@@
$\dot{X}_{1}$ $=\frac{1}{2\pi}\sin T\int_{-\infty}^{T}\int_{-\infty}^{\infty}\tilde{g}(K,0,\eta_{K}T^{\prime})\sum_{n=-\infty}^{\infty}\frac{nJ_{n}(K)}{K}e^{inT^{\prime}}\mathrm{d}K\mathrm{d}T^{\prime}$
$\quad\,-\frac{1}{2\pi}\cos T\int_{-\infty}^{T}\int_{-\infty}^{\infty}\tilde{g}(K,0,\eta_{K}T^{\prime})\sum_{n=-\infty}^{\infty}\frac{J_{n-1}(K)-J_{n+1}(K)}{2i}e^{inT^{\prime}}\mathrm{d}K\mathrm{d}T^{\prime}.$ (23)
@@ 3.1 Change in perpendicular energy
$\frac{v_{\perp}^{2}}{v_{\perp 0}^{2}}=1+2\varepsilon\left(\dot{X}_{0}\dot{X}_{1}+\dot{Y}_{0}\dot{Y_{1}}\right)_{T\to\infty}+O(\varepsilon^{2}).$ (25)
@@
$\Delta\sim 2i\int_{-\infty}^{\infty}\sum_{n=-\infty}^{\infty}\textrm{sgn}(n)\,\textrm{Res}[A(n,K,s),s_{*}]\exp\left(-\frac{|n\mathrm{Im}\left\{s_{*}\right\}|}{\eta_{K}}\right)\mathrm{d}K.$ (28)
@@ 3.3 Scattering contours
Let us for the moment assume that the electromagnetic fluctuations are from a propagating wave or superposition of waves, with a parallel phase velocity $v_{\mathrm{ph}}(K)$, so that $\tilde{b}=\tilde{b}(K,Z-(v_{\mathrm{ph}}(K)/v_{\perp 0})T)$, and $\tilde{g}=\tilde{g}(K,Z-(v_{\mathrm{ph}}/v_{\perp 0})T)$. The results derived in the previous sections do not require this, but it will allow us to make contact with the usual quasilinear theory of cyclotron heating. It may also be directly applicable to the turbulence in the solar wind, which can be highly imbalanced: dominated by outward-going Alfvén waves, for which $v_{\mathrm{ph}}=v_{\mathrm{A}}$. More generally, this situation could potentially also apply to nonlinear solitary waves which have a single effective $v_{\mathrm{ph}}$ due to the nonlinearity balancing dispersion (Kakutani and Ono, 1969; Kawahara, 1969; Hasegawa and Mima, 1976; Mjølhus and Wyller, 1986; Mallet, 2023). We note that we are ignoring the possibility of purely waves with phase velocity only in the perpendicular direction, and also electrostatic waves (which to have parallel phase velocities must also have $E_{z}\neq 0$ so that Faraday’s law is satified).
@@ 4 Low-$\beta$ Alfvénic turbulence
In the Alfvénic turbulence present in the solar wind and corona, both $\varepsilon$ and $\eta$ depend on $k_{\perp}$. We will assume a relatively low $\beta$, so that the kinetic reduced electron heating model (KREHM) equations (Zocco and Schekochihin, 2011) may be used; we also assume that while $k_{\perp}\rho_{p}\sim k_{\perp}\rho_{s}\sim 1$, $k_{\perp}d_{e}\ll 1$, so the electrons are isothermal; where the thermal proton gyroradius $\rho_{p}=v_{th\rm{p}}/\Omega_{p}$ (different from $\rho$!), the ion sound radius is $\rho_{s}=\sqrt{ZT_{e}/m_{i}}/\Omega_{p}$, with $\Omega_{p}=eB/m_{p}c$ the proton gyrofrequency, and $d_{e}=c/\omega_{pe}$ is the electron inertial length. We want to express our results in terms of what is experimentally observable in the solar wind; namely, the $\delta B_{\perp}$ fluctuation amplitude as a function of the perpendicular wavenumber $k_{\perp}$; we will write this in velocity units as $\delta b_{k}=\delta B_{k}/\sqrt{4\pi n_{0p}m_{p}}$. For now, we will also neglect intermittency in the turbulent fluctuation amplitude, supposing that we may characterise the amplitude at each scale by a single value $\delta b_{k}$. (The effects of intermittency will be examined in Sec. 4.4.) Moreover, we assume the critical balance, such that $\omega\sim k_{\perp}\delta u_{e}$, where the effective electron bulk flow velocity is (Zocco and Schekochihin, 2011)
@@
This is true not only for linear Alfvén waves, but statistically even in strongly nonlinear kinetic-Alfvén turbulence (Grošelj et al., 2018): in a similar sense to the fact that $\delta u\sim\delta b$ in strong MHD turbulence (Maron and Goldreich, 2001). Since $\Gamma_{0}\approx 1-k_{\perp}^{2}\rho_{p}^{2}/2$ for $k_{\perp}\rho_{p}\ll 1$, but $\Gamma_{0}\approx 0$ for $k_{\perp}\rho_{p}\gg 1$, $\alpha_{k}\to 1$ as $k_{\perp}\rho_{p}\ll 1$ (Alfvén waves) and $\alpha_{k}\to k_{\perp}\rho_{\rm th}$ for $k_{\perp}\rho_{p}\gg 1$ (kinetic Alfvén waves). Thus, we have
@@
where we have neglected $k_{\parallel}v_{\parallel 0}\ll\omega\sim k_{\parallel}v_{\mathrm{A}}$ since $\beta_{i}$ is small, and as promised added an undetermined constant $c_{2}$ accounting for (several) prefactors of order unity we have neglected. This could in principle be corrected for the slowing down of the turbulent cascade due to dynamic alignment (Boldyrev, 2006; Chandran et al., 2015; Mallet and Schekochihin, 2017) and/or imbalance (Schekochihin, 2022; Chandran et al., 2025). A more important limitation is that the KREHM equations assume that $\omega/\Omega_{p}\ll 1$. We will rather flagrantly ignore this restriction in the following analysis: while it could be corrected for by using a more accurate dispersion relation, we believe it does not impact our results in a significant way. The parameter $\varepsilon_{k_{\perp}}=Ec/Bv_{\perp 0}$ controlling the amplitude of the electric fields is similarly found [^4] Note that due to the smallness of $\beta$ implied by our use of the KREHM equations, there is no contribution to $\phi$ from parallel magnetic field fluctuations $\delta b_{\parallel}$ as appears in the stochastic heating theory of Cerri et al. (2021).: using $E\sim k_{\perp}\phi$ and using Eqs. (49) and (51),
@@
For the moment assuming that the ion component of the plasma is mainly protons [^5] This assumption is in fact implicit in Eqs. 4953 earlier in this section: the minor ions have sufficiently low abundances that they do not affect the KAW dispersion relation nor polarization properties of the fluctuations., $v_{\rm th}=v_{th\rm p}$, $\rho_{\rm th}\sim\rho_{p}$, and $\Omega_{i}=\Omega_{p}$, we have
@@
We will return to the subject of minor ions in Sec. 4.5. Taking $k\rho_{p}\sim 1$, we recover the expression given in Chandran et al. (2010), where the exponential suppression factor depends only on the amplitude of the fluctuations: the equivalence between exponential suppression factors based on the timescale and the amplitude for $k\rho_{p}\sim 1$ was first noticed by Cerri et al. (2021). However, we can now assess the scale dependence of the heating directly. It is interesting to look at this in different limits. For $k_{\perp}\rho_{p}\ll 1$, we have $1-\Gamma_{0}\approx k_{\perp}^{2}\rho_{p}^{2}/2$, $\alpha_{k}\sim 1$, and $J_{1}(k_{\perp}\rho_{p})\sim k_{\perp}\rho_{p}/2$. Then,
@@
The only difference is in the exponential suppression factor; the speedup in the frequency for small-scale fluctuations means we get an extra factor of $k_{\perp}\rho_{p}$ there. Were $\delta b_{k}$ independent of scale, the heating rate becomes monotonically larger towards smaller scale, since the frequency increases. Eqs. (5658) can be compared with the heating rate (due to Landau/transit-time damping) in Alfvénic turbulence within the gyrokinetic model (Quataert, 1998; Howes, 2010), $Q_{\perp}\sim k_{\perp}\delta b_{k}^{3}\gamma_{p}/\omega$: since $\gamma_{p}/\omega$ is rather small at low $\beta$, magnetic moment breaking may be the dominant heating mechanism despite the exponential suppression factor. Future work will assess this statement in a more satisfactory manner.